Rogen Corporation manufactures a single product. The standard cost per unit of product is shown below.

Direct materials—1 pound plastic at $6 per pound $ 6.00
Direct labor—0.50 hours at $11.75 per hour 5.88
Variable manufacturing overhead 3.50
Fixed manufacturing overhead 1.50
Total standard cost per unit $16.88


The predetermined manufacturing overhead rate is $10 per direct labor hour ($5.00 ÷ 0.50). It was computed from a master manufacturing overhead budget based on normal production of 2,950 direct labor hours (5,900 units) for the month. The master budget showed total variable costs of $20,650 ($7.00 per hour) and total fixed overhead costs of $8,850 ($3.00 per hour). Actual costs for October in producing 3,600 units were as follows.

Direct materials (3,800 pounds) $ 23,180
Direct labor (1,670 hours) 20,291
Variable overhead 13,678
Fixed overhead 6,622
    Total manufacturing costs $63,771


The purchasing department buys the quantities of raw materials that are expected to be used in production each month. Raw materials inventories, therefore, can be ignored.

(a)

Compute all of the materials and labor variances. (Round answers to 0 decimal places, e.g. 125.)

Total materials variance $

FavorableNeither favorable nor unfavorableUnfavorable

Materials price variance $

FavorableNeither favorable nor unfavorableUnfavorable

Materials quantity variance $

Neither favorable nor unfavorableUnfavorableFavorable

Total labor variance $

Neither favorable nor unfavorableFavorableUnfavorable

Labor price variance $

FavorableNeither favorable nor unfavorableUnfavorable

Labor quantity variance $

FavorableUnfavorableNeither favorable nor unfavorable


(b)

Compute the total overhead variance.

Total overhead variance $

FavorableNeither favorable nor unfavorableUnfavorable
Answer #1

3600 X SQ = 1.00 = 3600 Particulars Direct Material SQ AQ SP АР 3600 6 3800 6.1 Total Material variance (TMV) = (SQ * SP) - (SH = 3600 X 0.50 = 1800 Particulars SH SR АНР AR Direct Labour 1800 11.75 1670 12.15 Labour Price Variance (LPV) = ( SR - AR